OpenStudy (anonymous):

DEMO FOR A PRESENTATION PLEASE DON"T ANSWER What is x? -2x+32x=24/3

OpenStudy (anonymous):

-2x+32x=24/3 First Divide 24/3 -2x+32x=8 Then add the "x" numbers together 30x=8 Lastly, divide 30 out of both sides 30x/30=8x/30 x=0.26 repeating or 0.27

OpenStudy (anonymous):